# Deep Reinforcement Learning for Controlled Traversing of the Attractor Landscape of Boolean Models in the Context of Cellular Reprogramming

### v2

@@ 5.1 Pseudo-attractor states identification procedure
In light of Def. 4.1 and Obs. 4.3, the associated pseudo-attractor of an attractor of size $20$ can be identified only if the stationary distribution on the attractor is uniform. If the attractor size is less than 20, it is still possible to include all attractor states in the pseudo-attractor even if the distribution is non-uniform. Notice that with decreasing size of the unknown attractor, our procedure allows more and more pronounced deviations from the uniform distribution while preserving the complete attractor detection capability provided the stationary probabilities of all attractor states are above the threshold.
If an attractor is of size larger than $20$ states, Step I of our procedure with $5\%$ identification threshold will identify the associated pseudo-attractor only if the stationary distribution is non-uniform and the pseudo-attractor will contain only the most frequently revisited states. The maximum possible size of the identified pseudo-attractor in this case is $19$, which follows from Obs. 5.1. This is a desired property of our procedure as it keeps the number of pseudo-attractor states manageable which has significant positive influence on stabilising the model training as will be discussed below.
The environment pre-processing phase provides an initial set of pseudo-attractor states. The initial set is expanded in Step II during the model training phase. Step II-1 allows to identify plausible fix-point attractors. Step II-2 enables the identification of plausible multi-state attractors. However, here the focus is on smaller attractors than in the case of Step 1: we classify states as pseudo-attractor states if they are revisited at least $15\%$ of time, which corresponds to attractors of size 6. This is to restrict the number of spurious pseudo-attractor states in order to stabilise model training as explained next.
We have encoutered an issue related to late discovery of pseudo-attractor states during training. As can be observed in Fig. 1(a), the procedure may detect a new pseudo-attractor at any point in time which destabilises training: a new state is detected at around 90000 steps, which causes abrupt, significant increase of the average episode length. We propose a remedy to this problem in Sec. 5.2. Our experiments with small networks, i.e., ones for which exact attractors could be computed, revealed that it is beneficial to underestimate the set of attractor states in Step I of the procedure as the missed ones are usually discovered later during the training phase.
For big networks, e.g., with hundreds of nodes, the set of pseudo-attractors may take a long time to stabilise. Yet this approach provides us with the ability to process networks too big to be handled by traditional methods. The computations of pseudo-attractors can be parallelised in a rather straightforward way to speed up the detection. Furthermore, the notion of pseudo-attractors can easily be generalised to other types of GRN models, e.g., PBNs with perturbation, which is yet another well-established GRN modelling framework.
5.2 Exploration probability boost
The approach of [14] implements the $\varepsilon$-greedy policy in order to balance exploitation and exploration of the DRL agent during training. The $\varepsilon$-greedy policy introduces the exploration probability $\varepsilon$ and with probability $1-\varepsilon$ follows the greedy policy, i.e., it selects the action $a^{*}=\arg\max_{a\in\mathcal{A}}Q(s,a)$, or with the $\varepsilon$ exploration probability selects a random action. We set the initial $\varepsilon$ value to 1 and linearly decrease it to 0.05 over the initial 3000 steps of training.
Combining the original $\varepsilon$-greedy policy with online pseudo-attractor states identification gives rise to unstable training. When trying to train the DRL agent for our control problem while keeping identifying pseudo-attractor states during training, stability issues discussed in Sec. 4.1 were observed. To alleviate this negative influence on training, we introduce the exploration probability boost (EPB) to the $\varepsilon$-greedy policy. The idea of EPB is to increase the exploration probability $\varepsilon$ after each discovery of a new pseudo-attractor to $\max(\varepsilon,0.3)$ if the current value of $\varepsilon$ is less than 0.3. After the increase, the linear decrease to 0.05 follows with the rate of the initial decrease. As revealed by our experiments, this simple technique makes learning much more stable. This is illustrated in Fig. 1(b), where the agent discovers new pseudo-attractor states at around the 150000-th training step and the use of the improved $\varepsilon$-greedy policy allowed us to reduce the increase of the average episode length in a significant way and resulted in a quick return to the previously trained low value of the average episode length.
((a)) Training without EPB.
((b)) Training with EPB.
Figure 2: Examples of average episode lengths during training run with and without EPB. New pseudo-attractor states are being identified during training.
5.3 pbn-STAC implementation
We implement pbn-STAC as a fork of gym-PBN [8], an environment for modelling of PBNs, and pbn-rl [9], a suite of DRL experiments for a different PBN control problem formulated in [14]. In pbn-STAC, we have adapted the original code of gym-PBN and pbn-rl to our formulation of the PBN control problem, i.e., the source-target attractor control. First, we extend gym-PBN by adding the asynchronous PBN environment to it. Second, to allow for simultaneous perturbation of a combination of genes within a DRL action, we replace the original DDQN architecture with the BDQ architecture [25], which, contrary to DDQN, scales linearly with the dimension of the action space. The architecture of our BDQ network is depicted in Fig. 3.
Figure 3: Schematic illustration of the BDQ network architecture
Third, we implement the pseudo-attractor states identification procedure and the exploration probability boost technique. Finally, the framework takes as input a source-target pair of attractors or pseudo-attractor states. In the case of a multi-state target attractor, a training episode is regarded as successful if any of the target attractor states is reached. For a multi-state source attractor, we uniformly sample one of its states and set it as the initial state. In this way, different source attractor states are considered as initial during DRL agent training. Our DRL-based framework for the source-target attrator control is made available via the dedicated pbn-STAC GitHub repository [28].
6 Experiments
6.1 BN and PBN models of GRNs
Melanoma models. We infer BN and PBN environments of various sizes for the melanoma GRN using the gene expression data provided by Bittner et al. in [6]. This is a well-known dataset on melanoma, which is extensively studied in the literature, see, e.g., [14, 5, 7]. To infer the BN/PBN structures, we follow the approach of [14] implemented in gym-PBN [14]. It is based on the coefficient of determination (COD), which is a measure of how well the dependent variable can be predicted by a model, a perceptron in the case of [14].
The original dataset of Bittner et al. is quantised by the method of [14]. Then, the BN and PBN models of sizes 7, 10, and 28 are obtained from these data. The models are denoted as BN-x or PBN-x, respectively, where x is the number of genes. To infer the predictors for the models, we set the number of predictors for each gene to 1 for BN models and to 3 for PBN models. For each gene, the algorithm selects the Boolean functions with the maximum COD values. For more details on the inference method, we refer to [14].
Case study of B. bronchiseptica. We test our DRL-based control framework on an existing model of a real biological system, i.e., the network of immune response against infection with the respiratory bacterium Bordetella bronchiseptica, which was originally proposed and verified against empirical findings in [26]. The computational model, denoted IRBB-33, is an asynchronous BN consisting of 33 genes.
6.2 Performance evaluation methodology
We evaluate the performance of pbn-STAC in solving the control problem formulated in Sec. 4 on BN and PBN models of melanoma of various sizes, i.e., incorporating 7, 10, and 28 genes. Moreover, we consider IRBB-33, the 33-genes BN model. The dynamics under the asynchronous update mode is considered for all models. The evaluation consists of the agent interacting with the environment by taking actions, where an action consists of flipping the values of a particular subset of genes in an attractor or pseudo-attractor state. We recover a control strategy for a given source-target pair learned by a trained DRL agent byinitialising the BN/PBN environment with the source and target and letting it run while applying the actions suggested by the DRL agent in the source and all the intermediate (pseudo-)attractor states encountered on the path from the source to the target. To evaluate the performance of pbn-STAC on a particular BN/PBN model, we recover control strategies for all possible ordered source-target pairs of the model’s attractors or pseudo-attractor states. For all the BN models of melanoma and IRBB-33, we are able to compute all their attractors and optimal control strategies for all pairs of attractors using the CABEAN software tool with the attractor-based sequential instantaneous source-target control (ASI) method. We use the information on the exact attractors and optimal control strategies for BN models as ground truth for the evaluation of pbn-STAC.
For PBN-7 and PBN-10, we compute the attractors with the NetworkX package [12], which facilitates the analysis of complex networks in Python. Unfortunately, due to very large memory requirements, we are unable to obtain the attractors of the 28-genes PBN model of melanoma with this approach, so we consider pseudo-attractor states. The optimal-length control strategies are obtained for PBN-7 and PBN-10 models by exhaustive search.
Notice that due to the nondeterministic nature of our environments, i.e., the asynchronous update mode, the results may vary between runs. Therefore, for each source-target pair, we repeat the run 10 times. For each recovered control strategy, we count its length, and record the information whether the target attractor is reached. For a given BN/PBN model, we report the percentage of successful control strategies found and the average length of the successful control strategies.
7 Results
7.1 Identification of pseudo-attractor states
We evaluate the performance of PASIP proposed in Sec. 5.1. For this purpose, we run pbn-STAC with PASIP on the considered BN and PBN models. We present the obtained results in Tab. 1. For each model, except the melanoma PBN-28 for which the exact attractors could not be obtained, we provide the information on the number of exact attractors, the total number of attractor states, and the total numbers of identified pseudo-attractor states with our procedure. We measure the precision of our approach defined as $\textrm{TP}/(\textrm{TP}+\textrm{FP})$, where TP is the number of true positives, i.e., the number of pseudo-attractor states that are attractor states, and FP is the number of false positives, i.e., the number of states identified as pseudo-attractor states which are not part of any of the network’s attractor. We can conclude that for all cases in which the exact attractors are known, our procedure does not introduce any FPs. Moreover, it can identify the attractor states with $100\%$ precision in all but one case, i.e., the PBN-28 network which has 2412 fix-point attractors and for which our procedure correctly identifies 1053 of them. This justifies our strong belief that running our procedure for longer time would result in definitely higher precision also in the case of BN-28. In summary, the presented results show that PASIP is reliable.
Model #Attr. #Attr. states #PA-states Precision
BN-7 6 6 6 100%
BN-10 26 26 26 100%
BN-28 2412 2412 1053 43.65%
IRBB-33 3 3 3 100%
PBN-7 4 4 4 100%
PBN-10 6 6 6 100%
PBN-28 unknown unknown 14 N/A
Table 1: Comparison of the number of exact attractor states and pseudo-attractor states identified by PASIP for various BN and PBN models. The fact that we were unable to obtain the exact attractors for the PBN-28 model is indicated with ‘unknown’. Attr. is short for Attracor and PA stands for Pseudo-attractor.
Model #Attractors Optimal Strategy pbn-STAC
BN-7 6 1.0 3.98
BN-10 26 1.1 2.14
BN-28 2412 1.1 -
IRBB-33 3 1.0 9.2
PBN-7 4 1.1 5.5
PBN-10 6 1.2 15.2
PBN-28 unknown unknown 60.7
Table 2: Average lengths of pbn-STAC control strategies and optimal control strategies obtained with CABEAN (BNs) or exhaustive search (PBNs) for all source-target pairs of individual models. The fact that we were unable to obtain the optimal strategy for the PBN-28 model is indicated with ‘unknown’.
7.2 Control of BN models of melanoma
We evaluate the ability of pbn-STAC to solve the control problem by comparing the obtained results to the optimal ASI control strategies computed with CABEAN. As can be seen in Tab. 2, the strategies obtained with pbn-STAC for larger BN models tend to be longer on average compared to the optimal ones. However the overhead is rather stable across different models. We investigate the issue of longer control strategies further by computing a histogram of control strategy lengths for the BN-7 model provided in Fig. 4(a). It is apparent that in most cases the control strategies are short and close to the optimal ones. Nevertheless, there are a few cases of longer control strategies that give rise to the higher average values. The longer strategies are present due to the fact that the interventions suggested by the trained DRL-agent often place the system in a so-called weak basin of attraction of an attractor, i.e., a set of states from which the attractor is reachable, but not necessarily – the dynamics can still lead the system to another attractor from these states due to non-determinism arising from the asynchronous update mode. The strategies computed by CABEAN are optimal since they are obtained by considering the so-called strong basins of attraction, i.e., states from which only a single attractor can be reached. Nevertheless, determining strong basins is challenging, not to say impossible, for large networks (see [18] for details). In light of this and the fact that pbn-STAC can handle larger networks, the obtained results can be seen as reasonable and acceptable.
Unfortunately, due to the huge number of attractors of the BN-28 model, the training of pbn-STAC on this model needs to be run for much longer time and we did not manage to finish it within our time limits. Notice that our training procedure considers all ordered pairs of the attractors. Handling such cases requires further research.
((a)) Usual reward
((b)) Improved reward
7.3 Control of the IRBB-33 model
In the case of the IRBB-33 network, we have to modify the reward scheme. As can be seen in Fig. 3(a), the reward scheme introduced in Sec. 5, referred to as the mixed reward, does not lead to any improvement of the average episode length during training over 200 000 training steps. After trying different reward schemes for this network (data not shown), we found that the following scheme
$R_{a}(s,s^{\prime})=-|a|+100*(\mathbbm{1}_{TA}(s^{\prime})-1),$
It improves the training of the DRL agent significantly, as can be seen in Fig. 3(b), where the convergence is achieved in tens of thousands of steps.
((a)) BN-7
((b)) BN-10
((c)) PBN-28
((d)) BN-33
Figure 5: Histogram of the control strategy lengths for the BN-7 model.
The average control strategy length obtained with pbn-STAC is $9.2$, as presented in Tab. 2. The length is again larger than in the optimal case, but the overhead is stable with respect to the results obtained for the BN models of melanoma. Again, as can be observed in Fig. 4(d), in the majority of cases the strategies are of length one, which perfectly corresponds with the optimal strategy. Unfortunately, there are a few very long ones, which give rise to the higher average value.
7.4 Control of PBN models of melanoma
We run the pbn-STAC control framework on the three PBN models of melanoma. For PBN-28, we are not able to compute the set of exact attractors, but we identify 14 pseudo-attractor states. Unfortunately, we can not obtain the optimal control strategies for this network with exhaustive search.
As shown in Tab. 2, the control strategies found by pbn-STAC are on average longer than the optimal ones. Moreover, their lengths seem to increase with the size of the network faster than in the case of BN models. Unfortunately, the optimal result is not available for the PBN-28 model to make a comparison. Although the average length for the this network is high, the distribution is heavily skewed with a long tail of longer control strategies as can be seen in Fig. 4(c). Nevertheless, once again the majority of the source-target pairs are controllable with very few interventions. This characteristic of the control strategies obtained with pbn-STAC is consistent across different models and theirs types.
8 Conclusions
In this study we formulated a control problem for the BN and PBN frameworks under the asynchronous update mode that corresponds to the problem of identifying effective cellular reprogramming strategies. We have developed and implemented a computational framework, i.e., pbn-STAC, based on DRL that solves the control problem. It allows to find proper control strategies that drive a network from the source to the target attractor by intervening only in other attractor states that correspond to phenotypical functional cellular states that can be observed in the lab. Since identifying attractors of large BNs/PBNs is a challenging problem by itself and we consider our framework as a contribution towards developing scalable control methods for large networks, we introduced the notion of a pseudo-attractor and developed a procedure that identifies pseudo-attractor states during DRL agent training. We evaluate the performance of pbn-SEC on a number of networks of various sizes and a biological case study and compare its solutions with the exact, optimal ones wherever possible.
The obtained results show the potential of the framework in terms of effectiveness and at the same time reveal some bottlenecks that need to be overcome to improve the performance. The major identified issue is related to the long tails of the distributions of the lengths of strategies identified by pbn-SEC, i.e., there are many strategies of lengths close to the optimal ones, and a few which are very long. This negatively influences the average value. Addressing this problem would allow us to significantly improve the performance of pbn-STAC and make it effective on large models. We consider these developments and the evaluations of pbn-STAC on models of large sizes as future work.
Finally, we perceive our framework as rather straightforwardly adaptable to other types of PBNs, such as PBNs with perturbations, or Probabilistic Boolean Control Networks.

### v3

@@ Appendix C Size of a pseudo-attractor
Since all states of $A$ are positive recurrent in the PBN underlying Markov chain restricted to $A$, it holds that $\mathbb{P}_{A}(s)>0$ for all $s\in A$. We start with the first case, where 100 mod $0100\text{ mod }k=$ 0 and $|A|>\frac{100}{k}$. A state $s$ is identified as a pseudo-attractor state by Step I if $\mathbb{P}_{A}(s)\geq\frac{k}{100}$. A pseudo-attractor associated with $A$ will be of maximum size whenever the stationary distribution maximises the number of states with $\frac{k}{100}$ probability. The maximum must be less than $\frac{100}{k}$. To show this, let $M$ denote the subset of states of $A$ with stationary distribution probability of $\frac{k}{100}$ and let assume that $|M|\geq\frac{100}{k}$. Then, $\sum_{s\in A}\mathbb{P}_{A}(s)=\sum_{s\in M}\mathbb{P}_{A}(s)+\sum_{s\in A\setminus M}\mathbb{P}_{A}(s)\geq\frac{k}{100}\cdot\frac{100}{k}+\sum_{s\in A\setminus M}\mathbb{P}_{A}(s)>1$, where the last inequality follows from the fact that all the states in $A\setminus M$ have non-zero stationary probabilities. Therefore, $\mathbb{P}_{A}$ cannot be a probability distribution. Hence, $|M|\leq\frac{100}{k}-1$. The upper bound is reached, for example, by any stationary distribution $\mathbb{P}_{A}$ such that the for the $\frac{100}{k}-1$ states in $M$ the probabilities are equal to $\frac{k}{100}$ and $\sum_{s\in A\setminus M}\mathbb{P}_{A}(s)=\frac{k}{100}$. This concludes the first case.
In the other case, let $\epsilon:=\frac{100}{k}-\left\lfloor\frac{100}{k}\right\rfloor$. If $|A|\leq\left\lfloor\frac{100}{k}\right\rfloor$, then the inequality holds since $PA\subseteq A$ by definition. The upper bound is reached if $|A|=\left\lfloor\frac{100}{k}\right\rfloor$ and, for example, the stationary distribution probabilities for $|A|-1$ states in $A$ are equal to $\frac{k}{100}$ and for the remaining state the probability is $\frac{k}{100}\cdot\left(1+\epsilon\right)$. If $|A|>\left\lfloor\frac{100}{k}\right\rfloor$ in the second case, then $\epsilon>0$, i.e., $\epsilon\in(0,1)$. The maximum possible number of states of $A$ with stationary probability equal $\frac{k}{100}$ must be less than $\left\lfloor\frac{100}{k}\right\rfloor+1$. To see this, let $M$ be the set of such states of $A$ and let assume that $|M|\geq\left\lfloor\frac{100}{k}\right\rfloor+1$. Then, $\sum_{s\in A}\mathbb{P}_{A}(s)>\frac{k}{100}\cdot\left(\left\lfloor\frac{100}{k}\right\rfloor+1\right)>\frac{k}{100}\cdot\left(\left\lfloor\frac{100}{k}\right\rfloor+\epsilon\right)=1$. This contradicts the fact that $\mathbb{P}_{A}$ is a probability distribution on $A$. Finally, we proceed to show that the upper bound can be reached also in this scenario. This is the case, for example, for a stationary probability distribution $\mathbb{P}_{A}$ on $A$ where $\left\lfloor\frac{100}{k}\right\rfloor$ states of set $M$ have probabilities $\frac{k}{100}$ and all the states in $A\setminus M$ have non-zero probabilities such that $\sum_{s\in A\setminus M}\mathbb{P}_{A}(s)=\epsilon$. This is possible since $\epsilon>0$. ∎
Example 3.
For the network defined in the running Example 1, the true attractor states are $(0,0,0)$, $(1,1,0)$, and $(1,1,1)$. By running $100$ simulations of this network for $1000$ steps each, the following counts of revisits for the individual PBN states were obtained: [47852[47852[ 47852, $34724$, $17173$, $74$, $65$, $46$, $34$, 32]32]32 ]. Thus, the percentages of time spent in each state are as follows: [0.479[0.479[ 0.479, $0.347$, $0.172$, $0.074$, $0.065$, $0.046$, $0.034$, 0.032]0.032]0.032 ], with the four most visited states being $(0,0,0)$, $(1,1,1)$, $(1,1,0)$, and $(0,0,1)$, first three of which with over 15% revisiting frequency.
Appendix D Reinforcement learning
In this section, we provide a concise review of reinforcement learning topics relevant in the context of our study. For a comprehensive treatment of the subject, we refer to [25]. Reinforcement learning (RL) can be studied within a framework based on Markov decision processes. Herein, we focus on discrete-time and finite processes defined as follows.
Definition 6 (Markov decision process).
A discrete time, finite Markov decision process (MDP) is a quadruple $(\mathcal{S},\mathcal{A},P,R)$, where:
  • $\mathcal{S}$ is a finite set of states, referred to as the state space;
  • $\mathcal{A}$ is the action space, i.e., a finite set of actions that can be taken in each state;
  • $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the function of state-transition probabilities that determines the conditional probability $P_{a}(s,s^{\prime})=\textrm{Pr}(\{S_{t+1}=s^{\prime}\mid S_{t}=s,A_{t}=a\})$ of transitioning from state $s$ to $s^{\prime}$ after taking action $a$ for each $s,s^{\prime}\in\mathcal{S}$ and $a\in\mathcal{A}$;
  • $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function which determines the immediate reward $R_{a}(s,s^{\prime})$ received after transitioning from state $s$ to $s^{\prime}$ after taking action $a$ for each $s,s^{\prime}\in\mathcal{S}$ and $a\in\mathcal{A}$.
The index $t\in\mathbb{Z}_{+}$ denotes a time step from the sequence of discrete time steps $0,1,2,\ldots$. The discrete random variables $S_{t}$, $A_{t}$, and $R_{t}$ are the state, action, and immediate reward at time $t$, respectively.
The RL problem is meant to be a direct framing of the task of learning from interaction to achieve a goal. The decision-maker and learner is referred to as the agent. At each time step $t$, the agent interacts with the environment which comprises everything outside the agent by implementing a mapping $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ from states to probabilities of selecting an action $A_{t}$ in state $S_{t}$. This mapping is referred to as the agent’s policy, where $\pi(a\mid s)$ is the probability that $A_{t}=a$ given that $S_{t}=s$.
Within the MDP-based framework, the RL problem is formalised as the optimal control of MDPs. RL methods specify how the agent changes its policy as a result of its experiences with the goal of maximising the expected discounted return:
$G_{t}\triangleq\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1},$
where $0\leq\gamma\leq 1$ is the discount factor.
Nearly all RL algorithms involve estimating value functions. For MDPs, two such functions are commonly considered. First, the state-value function $v_{\pi}(s)$ defined as
$v_{\pi}(s)\triangleq\mathbb{E}_{\pi}[G_{t}\mid S_{t}=s]=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}\bigm{|}S_{t}=s\right]\!,$
where $\mathbb{E}_{\pi}[\cdot]$ is the expected value of a random variable given that the agent follows policy $\pi$. Second, the action-value function $q_{\pi}(s,a)$ is defined as the expected return starting from $s$, taking the action $a$, and thereafter following policy $\pi$:
$q_{\pi}(s,a)$ $\triangleq\mathbb{E}_{\pi}[G_{t}\mid S_{t}=s,A_{t}=a]$
$=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}\biggm{|}S_{t}=s,A_{t}=a\right]\!.$
The value functions satisfy particular recursive relationships expressed in the form of Bellman equations, i.e., one for $v_{\pi}$:
$v_{\pi}(s)=\sum_{a\in\mathcal{A}}\pi(s,a)\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\Bigl{[}R_{a}(s,s^{\prime})+\gamma v_{\pi}(s^{\prime})\Bigr{]}$
and one for $q_{\pi}$:
$q_{\pi}(s,a)=\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\Bigl{[}R_{a}(s,s^{\prime})+\gamma\sum_{a^{\prime}\in\mathcal{A}}\pi(s^{\prime},a^{\prime})q_{\pi}(s^{\prime},a^{\prime})\Bigr{]}.$
Value functions define a partial order over policies, i.e., $\pi\geq\pi^{\prime}$ if and only if $v_{\pi}(s)\geq v_{\pi}^{\prime}(s)$ for all $s\in\mathcal{S}$. There always exists at least one policy which is better than or equal to all other policies [25]. Such policy is referred to as an optimal policy and all optimal policies are denoted by $\pi_{*}$. Optimal policies share the same optimal state-value function, denoted $v_{*}$, and defined as
$v_{*}(s)\triangleq\max_{\pi}v_{\pi}(s),$
for all $s\in\mathcal{S}$. They also share the same optimal action-value function, denoted $q_{*}$, which is defined as
$q_{*}(s,a)\triangleq\max_{\pi}q_{\pi}(s,a),$
for all $s\in\mathcal{S}$. It holds that
$q_{*}(s,a)=\mathbb{E}_{\pi_{*}}\bigl{[}R_{t+1}+\gamma v_{*}(S_{t+1})\bigm{|}S_{t}=a,A_{t}=a\bigr{]}.$
The optimal value functions satisfy the Bellman optimality equations, i.e.,
$v_{*}(s)$ $=\max_{a\in\mathcal{A}}\mathbb{E}\bigl{[}R_{t+1}+\gamma v_{*}(S_{t+1})\bigm{|}S_{t}=s,A_{t}=a\bigr{]}$
$=\max_{a\in\mathcal{A}}\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\bigl{[}R_{a}(s,s^{\prime})+\gamma v_{*}(s^{\prime})\bigr{]}$
and
$q_{*}(s,a)$ $=\mathbb{E}\bigl{[}R_{t+1}+\gamma\max_{a^{\prime}\in\mathcal{A}}q_{*}(S_{t+1},a^{\prime})\bigm{|}S_{t}=s,A_{t}=a\bigr{]}$
$=\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\bigl{[}R_{a}(s,s^{\prime})+\gamma\max_{a^{\prime}\in\mathcal{A}}q_{*}(s^{\prime},a^{\prime})\bigr{]},$
where the expectations are taken without reference to any specific policy since the value functions are optimal.
For finite MDPs, the Bellman optimality equations have unique solutions independent of the policy. If the dynamics of the MDP in known, i.e., the $P$ and $R$ functions, one can solve the Bellman optimality equations. Once the optimal state-value function $v_{*}$ is known, any policy that is greedy with respect to $v_{*}$ is an optimal policy.
Optimal policies can be computed with dynamic programming (DP) algorithms, such as policy iteration or value iteration, given a perfect model of the environment in the form of an MDP. DP algorithms are obtained by turning Bellman equations into update rules that improve the approximations of the desired value functions. However, the utility of DP algorithms is limited due to the perfect model assumption and its computational expense. Therefore, for large state space environments or cases there the model of the environment is unavailable, Monte Carlo methods or temporal-difference learning (TD) methods are used, see [25].
D.1 Q-learning
Q-learning is a TD control algorithm and is one of the most popular model-free methods for solving RL problems. It is used to learn the optimal action-selection policy. The one-step Q-learning, which is the simplest form of Q-learning, is defined by the update rule
$Q(S_{t},A_{t})\leftarrow Q(S_{t},A_{t})+\alpha\bigl{[}R_{t+1}+\gamma\max_{a\in\mathcal{A}}Q(S_{t+1},a)-Q(S_{t},A_{t})\bigr{]}\!,$
where $0<\alpha\leq 1$ is the learning rate, which determines how fast past experiences are forgotten by the agent, and the learned action-value function $Q$ directly approximates $q_{*}$. The algorithm can be outlined as follows. First, initialise $Q(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$ and start the environment at state $S_{0}$. For time steps $t=0,1,2,\ldots$, choose action $A_{t}$ according to the $\epsilon$-greedy policy, i.e.,
$A_{t}\leftarrow\begin{cases}\textrm{argmax}_{a\in\mathcal{A}}Q(S_{t},a)&\textrm{with probability }1-\epsilon\\ \textrm{uniformy random action in }\mathcal{A}&\textrm{with probability }\epsilon,\\ \end{cases}$
and take action $A_{t}$ in the environment. In result, the state is changed from $S_{t}$ to $S_{t+1}\sim P_{a}(s,\cdot)$. Observe $S_{t+1}$ and $R_{t+1}$. Finally, revise the action-value function $Q$ at state-action $(S_{t},A_{t})$ in accordance with the above update rule.
The Q-learning algorithm iteratively updates the Q-values based on observed transitions, gradually refining its estimates of the optimal action-value function. All that is required to assure convergence to $q_{*}$ is that all $(s,a)\in\mathcal{S}\times\mathcal{A}$ pairs continue to be updated. Notice that the convergence is guaranteed independently of the policy being followed. This property renders Q-learning an off-policy TD control algorithm and enables the use of the $\epsilon$-greedy policy, which is a simple technique for handling the exploration-exploitation tradeoff, to ensure the convergence. Once learned, the Q-values are used to provide the agent with the optimal policy for selecting actions in each state.
Appendix E Deep reinforcement learning
The Q-learning algorithm has proved successful for controlling small-scale environments. However, in the case of large state-action spaces, Q-learning cannot learn all action values in all states separately and can fail to converge to an optimal solution. To address this issue, the idea was conceived to learn a parameterised value function $Q(a,s;\theta)$ by updating the parameters $\theta$ in accordance with the standard Q-learning-derived rule:
$\theta_{t+1}\leftarrow\theta_{t}+\alpha\Big{[}R_{t+1}+\gamma\max_{a}Q(S_{t+1},a;\theta_{t})$
$-Q(S_{t},A_{t};\theta_{t})\Big{]}\nabla_{\theta_{t}}Q(S_{t},A_{t};\theta_{t}),$
resembling stochastic gradient descent. It was shown that as the agent explores the environment, the parameterised value function converges to the true $Q$-function, see, e.g., [25].
The above idea is realised in the deep RL (DRL) framework [13], which combines classical RL with artificial neural network (ANN) function approximation. In DRL, a function approximator is trained to estimate the Q-values. The approximator is a deep Q-network agent (DQN), i.e., a multi-layer ANN which for a given observed state of the environment $s$ outputs a vector of action values $Q(s,\cdot;\theta)$, where $\theta$ are the parameters of DQN. The aim of DQN is to approximate $q_{*}$ by iteratively minimising a sequence of loss functions $L_{t}(\theta_{t})$:
$L_{t}(\theta_{t})=\left(Y_{t}^{\textrm{DQN}}-Q(S_{t},A_{t};\theta_{t})\right)^{2},$
where $Y_{t}^{\textrm{DQN}}$ is the DQN target defined as
$Y_{t}^{\textrm{DQN}}\triangleq R_{t+1}+\gamma\max_{a}Q(S_{t+1},a;\theta_{t}^{-}),$
while $\theta_{t}^{-}$ and $\theta_{t}$ are the parameters of the target and online networks, respectively. The target network is the same as the online network, but its parameter values are copied from the online network every $k$ steps and kept fixed at other steps.
E.1 Double Deep Q-Network
In the study of [15], the Double Deep Q-Network (DDQN) [28] was used to control PBNs. Both standard Q-learning and DQN uses the same values both to select and to evaluate an action, which can lead to selection based on overestimated values [28]. DDQN combines Double Q-learning [11] with Deep Q-learning [13] to address this issue and uses the following DDQN target
$Y_{t}^{\textrm{DDQN}}\triangleq R_{t+1}+\gamma Q(S_{t+1},\textrm{argmax}_{a}Q(S_{t+1},a;\theta_{t});\theta_{t}^{\prime}).$
The second set of parameters, $\theta_{t}^{\prime}$, enables fairly evaluation of the greedy policy. The parameters of the DDQN are updated by symmetrically switching the roles of $\theta_{t}$ and $\theta_{t}^{\prime}$.
E.2 Branching Dueling Q-Network
In our study, to allow for simultaneous perturbation of a combination of genes within a DRL action, we replace the original DDQN architecture of [15] with the Branching Dueling Q-Network (BDQ) architecture [26]. BDQs are designed to address complex and high-dimensional action spaces; they enhance the scalability and sample efficiency of RL algorithms in complex scenarios. BDQs can be thought of as an adaptation of the dueling network architecture [29] into the action branching architecture. A BDQ uses two separate ANNs, i.e., the target network for evaluation and the controller network for selection of actions. Furthermore, instead of using a single output layer for all actions, BDQ introduces multiple branches, each responsible for a specific subset of actions. This decomposition allows for parallel processing and shared representations, making it more efficient for the agent to learn and execute tasks with a diverse action space. BDQs avoid overestimating Q-values, can more rapidly identify action redundancies, and generalise more efficiently by learning a general Q-value that is shared across many similar actions. The BDQ architecture used in our study is depicted in Fig. 5.
Figure 5: Schematic representation of the BDQ architecture of the DRL agent in pbn-STAC.
Appendix F List of genes included in BN and PBN models of melanoma
The genes included in individual BN and PBN models of melanoma are shown in Tab. 3.
Table 3: Genes included in individual BN and PBN models of melanoma. The presence of a gene in a particular model is indicated with $+$ and its absence with --.
gene BN/PBN-7 BN/PBN-10 BN/PBN-30
reticulon 1 + + +
hydroxyacyl-Coenzyme A dehydrogenase/3-ketoacyl-Coenzyme A thiolase/enoyl-Coenzyme A hydratase (trifunctional protein), beta subunit + + +
ESTs + + +
pirin + + +
melan-A + + +
wingless-type MMTV integration site family, member 5A + + +
S100 calcium-binding protein, beta (neural) + + +
hexabrachion (tenascin C, cytotactin) - + +
CD63 antigen (melanoma 1 antigen) - + +
endothelin receptor type B - + +
annexin A2 - + +
phosphofructokinase, liver - - +
phosphoglycerate mutase 1 (brain) - - +
synuclein, alpha (non A4 component of amyloid precursor) - - +
phospholipase C, gamma 1 (formerly subtype 148) - - +
growth associated protein 43 - - +
AXL receptor tyrosine kinase - - +
syndecan 4 (amphiglycan, ryudocan) - - +
EphA2 - - +
CD20 antigen - - +
matrix metalloproteinase 3 (stromelysin 1, progelatinase) - - +
integrin, beta 1 (fibronectin receptor, beta polypeptide, antigen CD29 includes MDF2, MSK12) - - +
tropomyosin 1 (alpha) - - +
retinoid X receptor, alpha - - +
RAB2, member RAS oncogene family - - +
synuclein, alpha (non A4 component of amyloid precursor) - - +
Appendix G Results
Herein, we provide detailed information in the form of heatmaps and histograms on the control strategy lengths obtained for the considered BN models (Fig. 6 and Fig. 8) and PBN models (Fig. 7 and Fig. 9) discussed in the main text.
((a))
((b))
((c))
((d))
Figure 6: Heatmaps of strategy lengths for individual source-target (pseudo-)attractor states of BN models averaged over 10 runs.
((a))
((b))
((c))
Figure 7: Heatmaps of strategy lengths for individual source-target (pseudo-)attractor states of PBN models averaged over 10 runs.
((a))
((b))
((c))
((d))
Figure 8: Histograms of the number of successful control strategies vs their length for the BN models.
((a))
((b))
((c))
Figure 9: Histograms of the number of successful control strategies vs their length for the PBN models.